OpenStudy (anonymous):

(r+25).5 = (r-25)2 how do i solve this

OpenStudy (anonymous):

-58.33333333333333333333333333333

OpenStudy (anonymous):

ok you need a fraction

OpenStudy (anonymous):

oh by the way that is a .5

OpenStudy (anonymous):

Oh Im sorry Ill redo that part

OpenStudy (anonymous):

its okay i thought that something was wrong

OpenStudy (anonymous):

distribute (r+25).5=(r-25)2 .5(r)+5(25)= 2(r)-2(25) .5r+12.50=2r-50 Solve for r 5r+12.50=2r-50 +50 +50 ______________________ 5r+62.50=2r -.5r -.5r __________________ 62.50= -3r 62.50/-3= -3r/-3 ?= r

OpenStudy (anonymous):

neva mind i'll ask my teacher